# High-dimensional structure underlying individual differences in naturalistic visual experience

### v1

@@ 1 Introduction
The human brain transforms visual input into subjective perceptual experiences. When presented with identical visual stimuli, individuals exhibit idiosyncratic patterns of neural activity across cortical regions involved in visual processing  (?). Idiosyncratic neural activity patterns remain stable across time  (?, ?) and carry meaningful information about behavior  (?, ?). Individuals with shared personality traits  (?), shared context  (?), or shared perspectives  (?) have more similar neural responses during naturalistic narrative experience. These findings underscore the importance of understanding differences in neural representations underlying differences in perception and behavior across people. However, the full extent of variability within rich neural structure that gives rise to unique visual experience remains largely unknown.
@@
Here, we investigated the dimensionality of neural representational structure underlying individual differences. Specifically, we quantified the extent to which inter-individual variability can be characterized through latent codes spanning multiple dimensional ranges. In neural population coding, dimensionality refers to the minimum number of axes needed to specify each response’s position in representational space  (?). A prevailing view in cognitive neuroscience holds that cortical representations embed complex, high-dimensional stimuli into lower-dimensional latent subspaces to support generalization  (?, ?). However, recent studies leveraging large-scale neural datasets show that natural stimulus representations in visual cortex span at least thousands of dimensions, with dimensionality scaling unboundedly up to the limit of the data  (?, ?). This observation suggests that neural representations are more high-dimensional than previously appreciated. Crucially, dimensions beyond the leading, high-variance components encode stimulus information that may be relevant for perception but is not well-characterized by conventional analyses  (?). These findings raise a critical question: Are there reliable individual differences across many latent dimensions of neural responses during naturalistic visual experience, or are they concentrated in a subset of dimensions that explain most of the variance in neural data? If neural representations are organized across multiple scales of variance, then individual differences in these representations might similarly span many latent dimensions. Alternatively, although stimulus-related variance is embedded in high-dimensional subspaces, it is possible that only a low-dimensional subspace might be relevant for individual differences in visual experience  (?, ?, ?).
@@
In this work, we use cross-decomposition to examine individual differences across the full spectrum of dimensions underlying cortical population codes. Our method combines spectral analysis with cross-validated functional alignment to characterize how neural representations vary between individuals at multiple representational scales. Using fMRI data of subjects viewing naturalistic movies, we quantify both the reliability and behavioral significance of intersubject variability in movie-evoked neural responses throughout the visual cortical hierarchy. Our findings reveal that reliable individual differences in visual processing manifest as distinct coactivation patterns distributed across multiple scales of neural representation, from dominant patterns explaining large amounts of variance to subtler patterns explaining increasingly smaller amounts. Crucially, we find that intersubject variability patterns are qualitatively distinct at each dimensional range and cannot be explained by conventional intersubject correlation (ISC) measures, revealing unique aspects of individual neural processing distributed across multiple levels of granularity. These multidimensional neural response patterns predict individual differences in movie recall, with neural-behavioral correlations distributed over a wide range of representational dimensions. These results highlight the multiscale structure of individual differences and establish a new spectral framework that advances our understanding of the high-dimensional neural architecture underlying subjective visual experience.
@@ 2.1 High-dimensional structure in cortical responses to movies
One key to understanding neural population codes is characterizing how variance is distributed across latent dimensions of neural activity. Traditional approaches like principal component analysis (PCA) decompose neural responses into orthogonal dimensions sorted by variance. However, PCA captures both stimulus-driven and noise-related variance, potentially obscuring reliable stimulus-evoked patterns. Here, we employ cross-decomposition  (?) to isolate reliable stimulus-related variance in neural activity. Our approach isolates stimulus-dependent activations through cross-validation: basis sets derived from training data are used to compute covariance spectra on held-out test data. Like hyperalignment  (?), cross-decomposition performs the Procrustes transformation to map a pair of neural response data into a set of orthogonal latent dimensions that maximize the covariance between the data sets. Importantly, our method provides a spectral characterization of shared variance, decomposing the shared signal into latent dimensions ordered by explained covariance.
We hypothesized that stimulus-related variance in movie-evoked neural responses would be distributed across multiple orders of magnitude of latent dimensions. While previous work used cross-decomposition to estimate the dimensionality of neural representations evoked by natural images  (?), it has not been previously applied to dynamic, naturalistic movie responses. Here we applied cross-decomposition to a dataset of 43 subjects viewing four naturalistic movies in an fMRI scanner  (?). The dataset was originally curated to study individual differences in neural event segmentation during movie viewing and contained movie stimuli with highly varied content. Thus, these movies provide an excellent stimulus set for testing the generalization of our findings using held-out stimuli that are qualitatively distinct from the training stimuli. Specifically, we used a one-movie-out cross-validation scheme, which provided a stringent test of reliability. For a pair of subjects, we extracted shared latent dimensions from a training set of movies, with one movie held out for testing (Fig. 1a, left). During the testing phase, the held-out movie responses from both subjects were projected onto shared latent dimensions, yielding a cross-covariance spectrum for the held-out data (Fig. 1a, right). We averaged the cross-covariance spectra across movie-folds and repeated this procedure for every subject pair.
@@
Movie-evoked responses showed a characteristic power-law distribution of variance across latent dimensions in all examined regions, with cross-validated covariance spanning multiple orders of magnitude in each region (Fig. 1c). The smooth decay in covariance values across ranks showed that movie-evoked information is distributed across many dimensions. To verify that the covariance estimates over a wide range of ranks are statistically significant, we performed permutation tests by randomly block-shuffling movie-evoked responses across timepoints. We confirmed that up to  150 dimension ranks, the covariance statistics are significantly above the permuted spectra in all ROIs (Fig. S1). We thus focus on this range of dimensions for our analyses. We also examined the characteristic spatiotemporal scales of covariance patterns captured by each dimension and found a trend going from coarse- to fine-scale granularity across dimension ranks in both space and time (Fig. S2). Thus, the cross-decomposition analysis showed that the naturalistic movie viewing responses contained many dimensions of meaningful variance shared between individuals beyond the first few components, with successive dimensions reflecting finer patterns of cortical activity.
@@
Figure 1: Cross-decomposition analysis reveals many shared neural dimensions during movie viewing. (a) One-movie-out cross-validated cross-decomposition is used to identify reliable neural signal. Left: Shared dimensions from a subject pair are extracted from neural responses to three movies, with one movie held-out for testing. Right: The held-out movie responses are mapped onto these shared dimensions ranked by the covariance on the training set, generating a covariance spectrum that quantifies the amount of reliable signal on the test data. The procedure is repeated for each held-out movie to yield an average spectrum for each subject pair. (b) Regions of interest (ROIs) spanning the cortical hierarchy from low-level vision to high-level semantics: early visual areas (purple), ventral temporal (violet), lateral temporal (red), and posterior parietal-cingulate (orange) regions. (c) Cross-validated covariance spectra for each ROI. Violin plots show the distribution of reliable covariance across all subject pairs. The reliable stimulus-related covariance follows a power-law distribution across dimension-rank in all regions, demonstrating that individual neural response patterns span multiple orders of magnitude rather than being confined to low-dimensional subspaces.
@@ 2.2 Individual differences in high-dimensional cortical response patterns
To investigate this, we characterized patterns of individual differences and assessed their reliability across different ranges of dimensions (Fig. 2a). We captured subject-by-subject similarity patterns of shared information in neural responses using individual differences matrices (IDMs)  (?). We constructed IDMs for bins of latent dimensions for separate subsets of movies and computed the split-half reliability of individual differences across the full dimensional spectrum. If reliable individual differences span the full spectrum, this would indicate that the entire range of dimensions in these high-dimensional codes are potentially relevant to how individuals represent naturalistic stimuli. The reliability of individual differences in neural representations is important because it establishes the upper bound of the correlation between individual differences in neural codes and individual differences in behavior  (?).
@@
Figure 2: Reliable individual differences persist across multiple dimensions of shared neural representations during movie viewing. (a) Individual differences analysis compares patterns between independent sets of data to assess reliability. Covariance spectra are generated separately for neural responses to odd and even movies from each subject pair. These spectra are used to construct rank-wise individual differences matrices (IDMs), where red/blue colors indicate high/low dissimilarity between subjects (covariance values normalized for visualization purposes). Correlations between odd and even IDMs quantify reliability of individual difference patterns at each dimension rank, as shown in the hypothetical correlation plot on the right. (b) The observed correlations in all ROIs show significance across all examined dimension ranks up to two orders of magnitude. The persistence of significant correlations across high-dimensional ranks indicates that fine-grained neural response patterns in cross-validated covariance spectra contain reliable individual differences that cannot be reduced to simpler, coarse-scale properties detected by voxelwise intersubject correlation (ISC) measures. Violin plots show the bootstrap distribution of correlations between odd and even movie IDMs for each dimension rank bin, with 95% bootstrap confidence intervals (n=1,000 resamples). Dark shades represent correlations between original IDMs; lighter shades show correlations after controlling for coarse-scale effects through partialing out voxelwise ISC IDMs. Dots above each violin indicate statistical significance ($p<0.05$, Bonferroni-corrected).
@@
We found that individual differences were reliable across different subsets of movie stimuli throughout the entire spectrum of dimensions tested in all regions (Fig. 2b). The reliability of individual differences, quantified as correlations between IDMs derived from different sets of movies, remained significantly above zero at all dimensional ranges. This reliability was highest in low-rank dimensions but persisted even in high-rank dimensions, with statistically significant correlations extending to over a hundred dimensions in all examined regions (bootstrapped confidence intervals, n=1,000 resamples; $p<0.05$, Bonferroni-corrected).
@@
These surprisingly stable individual differences in high-dimensional codes are brought out by the cross-decomposition procedure, which increases the detectability of stimulus-related signal and reveals fine-grained idiosyncrasies in neural responses. Specifically, functional alignment transforms individual subjects’ data into a shared representational space where fine-grained, multivariate patterns of activity captured in latent dimensions can be directly compared across subjects. To rule out the possibility that these high-dimensional effects can be explained by coarse response properties, we computed individual differences based on average voxelwise similarity and partialed out these similarity patterns from our IDMs at each dimensional range (see Methods). This control analysis isolates individual differences encoded in the multivariate pattern structure beyond what can be explained by coarse-grained voxel correlations.
@@
While controlling for these coarse-scale effects modulated the magnitude of the IDM correlations, the overall pattern of reliable differences remained robust across the spectrum, and we continued to observe statistically significant reliability extending over a hundred dimensions in all regions (Fig. 2b). Thus, high-dimensional individual differences cannot be reduced to simpler, coarse-scale response properties but instead reflect differences in the fine-grained structure of neural representations.
@@
After establishing that individual differences are reliable across many ranks, we next asked: do different dimensional ranges capture similar or distinct patterns of individual differences? If each range of dimensions reflects unique aspects of individual neural processing, we would expect the correlation between IDMs from the same dimensional range to be significantly stronger than correlations between different ranges. To test this hypothesis, we computed correlations between IDMs derived from different dimensional ranges across different sets of movies. The resulting correlation matrices revealed a prominent diagonal structure across all cortical regions. The strongest correlations were observed between matching dimensional ranges, with the mean correlation along the diagonal significantly higher than off-diagonal elements (Fig. 3a). This pattern was particularly pronounced after controlling for coarse-grained response properties (Fig. 3b), revealing fine-grained patterns of variability that were obscured by coarse-scale effects. Furthermore, region-to-region comparisons showed that removing the coarse-grained patterns of intersubject correlations also brings out ROI-specific patterns of variability at each dimensional range (Fig. S3). These results indicate that each range of latent dimensions not only contains reliable individual differences but captures qualitatively distinct patterns of individual neural processing. This dimensional specificity reveals that individuals meaningfully differ on multiple scales of naturalistic stimulus processing.
@@
Figure 3: Distinct patterns of individual differences emerge across different dimensional ranges. (a) Correlation between individual difference matrices (IDMs) reveal stronger relationships between matching dimensional ranges than between different ranges. Each element represents the correlation between even-movie IDMs from one range (rows) and odd-movie IDMs from another range (columns). Diagonal values show same-range correlations that are consistently higher than off-diagonal values, indicating dimension-specific individual difference patterns. From left to right: early visual, ventral temporal, lateral visual, and posterior parietal-cingulate regions. (b) Same analysis after controlling for coarse-scale effects by partialing out voxelwise intersubject correlation (ISC) IDMs. The persistent diagonal structure indicates distinct, dimension-specific information in the residual, fine-grained patterns of individual differences. (c) Diagonal dominance (difference between on-diagonal correlation and mean off-diagonal correlation) remains significant across dimensional ranges before and after removing the coarse-grained effects. The diagonal dominance across ranks demonstrates that each dimensional range captures distinct aspects of individual neural processing during naturalistic viewing. Solid points show values from original matrices in a; lighter points show values after controlling for coarse-scale effects detected by conventional measures. All points shown are statistically significant (permutation tests, $p<0.05$).

### v2

@@ 1 Introduction
The human brain transforms visual input into subjective perceptual experiences. When presented with identical visual stimuli, individuals exhibit idiosyncratic patterns of neural activity across cortical regions involved in visual processing [Charest_Kriegeskorte_2015]. Idiosyncratic neural activity patterns remain stable across time [finn2015functional, gratton2018functional] and carry meaningful information about behavior [finn2015functional, Finn_Glerean_Khojandi_Nielson_Molfese_Handwerker_Bandettini_2020]. Individuals with shared personality traits [Finn_Corlett_Chen_Bandettini_Constable_2018], shared context [yeshurun2017same], or shared perspectives [lahnakoski2014synchronous] have more similar neural responses during naturalistic narrative experience. These findings underscore the importance of understanding differences in neural representations underlying differences in perception and behavior across people. However, the full extent of variability within rich neural structure that gives rise to unique visual experience remains largely unknown.
@@
Here, we investigated the dimensionality of neural representational structure underlying individual differences. Specifically, we quantified the extent to which inter-individual variability can be characterized through latent codes spanning multiple dimensional ranges. In neural population coding, dimensionality refers to the minimum number of axes needed to specify each response’s position in representational space [FUSI201666]. A prevailing view in cognitive neuroscience holds that cortical representations embed complex, high-dimensional stimuli into lower-dimensional latent subspaces to support generalization [Lehky_Kiani_Esteky_Tanaka_2014, dicarlo2007untangling]. However, recent studies leveraging large-scale neural datasets show that natural stimulus representations in visual cortex span at least thousands of dimensions, with dimensionality scaling unboundedly up to the limit of the data [stringer2019high, gauthaman2024]. This observation suggests that neural representations are more high-dimensional than previously appreciated. Crucially, dimensions beyond the leading, high-variance components encode stimulus information that may be relevant for perception but is not well-characterized by conventional analyses [gauthaman2024]. These findings raise a critical question: Are there reliable individual differences across many latent dimensions of neural responses during naturalistic visual experience, or are they concentrated in a subset of dimensions that explain most of the variance in neural data? If neural representations are organized across multiple scales of variance, then individual differences in these representations might similarly span many latent dimensions. Alternatively, although stimulus-related variance is embedded in high-dimensional subspaces, it is possible that only a low-dimensional subspace might be relevant for individual differences in visual experience [shine2019human, misra2021learning, lee2023hyper].
@@
In this work, we use cross-decomposition to examine individual differences across the full spectrum of dimensions underlying cortical population codes. Our method combines spectral analysis with cross-validated functional alignment to characterize how neural representations vary between individuals at multiple representational scales. Using fMRI data of subjects viewing naturalistic movies, we quantify both the reliability and behavioral significance of intersubject variability in movie-evoked neural responses throughout the visual cortical hierarchy. Our findings reveal that reliable individual differences in visual processing manifest as distinct coactivation patterns distributed across multiple scales of neural representation, from dominant patterns explaining large amounts of variance to subtler patterns explaining increasingly smaller amounts. Crucially, we find that intersubject variability patterns are qualitatively distinct at each dimensional range and cannot be explained by conventional intersubject correlation (ISC) measures [hasson2004intersubject], revealing unique aspects of individual neural processing distributed across multiple levels of granularity. These multidimensional neural response patterns predict individual differences in movie recall, with neural-behavioral correlations distributed over a wide range of representational dimensions. These results highlight the multiscale structure of individual differences and establish a new spectral framework that advances our understanding of the high-dimensional neural architecture underlying subjective visual experience.
@@ 2.1 High-dimensional structure in cortical responses to movies
One key to understanding neural population codes is characterizing how variance is distributed across latent dimensions of neural activity. Traditional approaches like principal component analysis (PCA) decompose neural responses into orthogonal dimensions sorted by variance. However, PCA captures both stimulus-driven and noise-related variance, potentially obscuring reliable stimulus-evoked patterns. Here, we employ cross-decomposition [gauthaman2024] to isolate reliable stimulus-related variance in neural activity. Our approach isolates stimulus-dependent activations through cross-validation: basis sets derived from training data are used to compute covariance spectra on held-out test data. Like hyperalignment [haxby2011common, guntupalli2016model], cross-decomposition performs the Procrustes transformation to map a pair of neural response data into a set of orthogonal latent dimensions that maximize the covariance between the data sets. We then obtain cross-validated covariance estimates for these aligned dimensions. Importantly, our method characterizes how information is distributed across the full spectrum of dimensions, allowing us to reveal the contribution of many subtle higher-rank dimensions that are typically overlooked.
We hypothesized that stimulus-related variance in movie-evoked neural responses would be distributed across orders of magnitude of latent dimensions. While previous work used cross-decomposition to estimate the dimensionality of neural representations evoked by natural images [gauthaman2024], it has not been previously applied to dynamic, naturalistic movie responses. Here we applied cross-decomposition to a dataset of 43 subjects viewing four naturalistic movies in an fMRI scanner [savasegal]. The dataset was originally curated to study individual differences in neural event segmentation during movie viewing and contained movie stimuli with highly varied content. Thus, these movies provide an excellent stimulus set for testing the generalization of our findings using held-out stimuli that are qualitatively distinct from the training stimuli. Specifically, we used a one-movie-out cross-validation scheme, which provided a stringent test of reliability.
Cross-decomposition requires two independent neural response datasets to identical stimuli to isolate stimulus-driven variance from noise. Although the movies were not shown repeatedly, all subjects viewed the same movies, allowing us to perform cross-decomposition using the responses from pairs of subjects. For a pair of subjects, we extracted shared latent dimensions from a training set of movies, with one movie held out for testing (Fig. 1A, left). During the testing phase, the held-out movie responses from both subjects were projected onto shared latent dimensions, yielding a cross-covariance spectrum for the held-out data (Fig. 1A, right). We averaged the cross-covariance spectra across movie-folds and repeated this procedure for every subject pair.
@@
Movie-evoked responses showed a characteristic power-law distribution of variance across latent dimensions in all examined regions, with cross-validated covariance spanning orders of magnitude of dimension ranks (Fig. 1C). Dimension ranks indicate the ordering of neural patterns from those explaining the most covariance (rank 1) to those explaining the least. To verify that the covariance estimates over a wide range of ranks are statistically significant, we performed permutation tests by randomly block-shuffling movie-evoked responses across timepoints (See Methods). We confirmed that up to approximately 150 dimension ranks, the covariance statistics are significantly above permuted spectra in all ROIs. We thus focus on this range of dimensions for our analyses We also examined the characteristic spatiotemporal scales of covariance patterns captured by each dimension and found a trend going from coarse- to fine-scale granularity across dimension ranks in both space and time (Fig. S1). Thus, the cross-decomposition analysis showed that the naturalistic movie viewing responses contained many dimensions of meaningful variance shared between individuals beyond the first few components, with successive dimensions reflecting finer patterns of cortical activity.
@@
Figure 1: Cross-decomposition analysis reveals many shared neural dimensions during movie viewing. (A) One-movie-out cross-validated cross-decomposition is used to identify reliable neural signal. Left: Shared dimensions from a subject pair are extracted from neural responses to three movies, with one movie held-out for testing. Right: The held-out movie responses are mapped onto these shared dimensions ranked by the covariance on the training set, generating a covariance spectrum that quantifies the amount of reliable signal on the test data. The procedure is repeated for each held-out movie to yield an average spectrum for each subject pair. (B) Regions of interest (ROIs) spanning the cortical hierarchy from low-level vision to high-level semantics: early visual areas (purple), ventral temporal (violet), lateral temporal (red), and posterior parietal-cingulate (orange) regions. (C) Cross-validated covariance spectra for each ROI. Violin plots show the distribution of reliable covariance across all subject pairs. The reliable stimulus-related covariance follows a power-law distribution across dimension-rank in all regions, demonstrating that individual neural response patterns span multiple orders of magnitude rather than being confined to low-dimensional subspaces. Dots above each violin indicate statistical significance ($p<0.05$, Bonferroni-corrected). Significance was tested against two baseline distributions: darker dots indicate significance against the randomly block-shuffled baseline, and lighter dots indicate significance against the random rotation baseline. (D) Example null spectra for early visual cortex. These plots show sample traces of individual null spectra (grey traces) and the mean of null spectra (dark line). These plots illustrate that individual null spectra fluctuate around zero, and they show that in expectation, the null distributions are flat. Top: mean permutation null spectra are centered at zero, indicating that stimulus desynchronization eliminates the heavy-tailed covariance structure. Bottom: mean random rotation spectra are similarly flat and centered at zero, confirming that arbitrary rotation does not lead to the observed covariance pattern.
@@ 2.2 Individual differences in high-dimensional cortical response patterns
To investigate this, we characterized patterns of individual differences and assessed their reliability across different ranges of dimensions (Fig. 2A). We captured subject-by-subject similarity patterns of shared information in neural responses using individual differences matrices (IDMs) [feilong2018reliable]. We constructed IDMs for bins of latent dimensions for separate subsets of movies and computed the split-half reliability of individual differences across the full dimensional spectrum. If reliable individual differences span the full spectrum, this would indicate that the entire range of dimensions in these high-dimensional codes are potentially relevant to how individuals represent naturalistic stimuli. The reliability of individual differences in neural representations is important because it establishes the upper bound of the correlation between individual differences in neural codes and individual differences in behavior [spearman1961proof].
@@
Figure 2: Reliable individual differences persist across multiple dimensions of shared neural representations during movie viewing. (a) Individual differences analysis compares patterns between independent sets of data to assess reliability. Covariance spectra are generated separately for neural responses to odd and even movies from each subject pair. These spectra are used to construct rank-wise individual differences matrices (IDMs), where red/blue colors indicate high/low dissimilarity between subjects (covariance values normalized for visualization purposes). Correlations between odd and even IDMs quantify reliability of individual difference patterns at each dimension rank, as shown in the hypothetical correlation plot on the right. (b) The observed correlations in all ROIs show significance across dimension ranks up to two orders of magnitude. The persistence of significant correlations across high-dimensional ranks indicates that fine-grained neural response patterns in cross-validated covariance spectra contain reliable individual differences that cannot be reduced to simpler, coarse-scale properties detected by voxelwise intersubject correlation (ISC) measures. Violin plots show the bootstrap distribution of correlations between odd and even movie IDMs for each dimension rank bin, with 95% bootstrap confidence intervals (n=1,000 resamples). Dark shades represent correlations between original IDMs; lighter shades show correlations after controlling for coarse-scale effects through partialing out voxelwise ISC IDMs. Grey horizontal line indicates zero correlation. Dots above each violin indicate statistical significance ($p<0.05$, Bonferroni-corrected).
@@
We found that individual differences were reliable across different subsets of movie stimuli throughout the entire spectrum of dimensions tested. The reliability of individual differences, quantified as correlations between IDMs derived from different sets of movies, remained significantly above zero at all dimensional ranges (bootstrapped confidence intervals, n=1,000 resamples; $p<0.05$, Bonferroni-corrected). This reliability was highest in low-rank dimensions but persisted even in high-rank dimensions. We additionally performed a whole-brain analysis that confirmed the strongest high-dimensional IDM correlations are in the posterior ROIs analyzed here (Fig. S4).
@@
These surprisingly stable individual differences in high-dimensional codes are brought out by the cross-decomposition procedure, which increases the detectability of stimulus-related signal and reveals fine-grained idiosyncrasies in neural responses. Specifically, functional alignment transforms individual subjects’ data into a shared representational space where fine-grained, multivariate patterns of activity captured in latent dimensions can be directly compared across subjects [haxby2011common]. To rule out the possibility that these high-dimensional effects can be explained by coarse response properties, we computed individual differences based on average voxelwise similarity and partialed out these similarity patterns from our IDMs at each dimensional range (see Methods). This control analysis isolates individual differences encoded in the multivariate pattern structure beyond what can be explained by coarse-grained voxel correlations.
@@
While controlling for these coarse-scale effects modulated the magnitude of the IDM correlations, the overall pattern of reliable differences remained robust across the spectrum, and we continued to observe statistically significant reliability extending over a hundred dimensions in all regions (Fig. 2B). We performed an additional validation that successively partials out correlations from preceding dimensions and confirmed that high-rank dimensions contain reliable information not captured by lower ranks (Fig. S2). Thus, high-dimensional individual differences cannot be reduced to simpler, coarse-scale response properties but instead reflect differences in the fine-grained structure of neural representations.
@@
After establishing that individual differences are reliable across many ranks, we next asked: do different dimensional ranges capture similar or distinct patterns of individual differences? If each range of dimensions reflects unique aspects of individual neural processing, we would expect the correlation between IDMs from the same dimensional range to be significantly stronger than correlations between different ranges. To test this hypothesis, we computed correlations between IDMs derived from different dimensional ranges across different sets of movies. We normalized each correlation by the geometric mean of the within-range reliabilities. Since dimensional bins above rank 100 did not show significant reliability, we excluded them from these analyses. The resulting correlation matrices revealed a prominent diagonal structure across all cortical regions. The strongest correlations were observed between matching dimensional ranges, with the mean correlation along the diagonal significantly higher than off-diagonal elements (Fig. 3A). This pattern was particularly pronounced after controlling for coarse-grained response properties (Fig. 3C), revealing fine-grained patterns of variability that were obscured by coarse-scale effects. Furthermore, region-to-region comparisons showed that removing the coarse-grained patterns of intersubject correlations also brings out ROI-specific patterns of variability at each dimensional range (Fig. S3). These results indicate that each range of latent dimensions not only contains reliable individual differences but captures qualitatively distinct patterns of individual neural processing. This dimensional specificity reveals that individuals meaningfully differ on multiple scales of naturalistic stimulus processing.
@@
Figure 3: Distinct patterns of individual differences emerge across different dimensional ranges. (a) Correlation between individual difference matrices (IDMs) reveal stronger relationships between matching dimensional ranges than between different ranges. Each element represents the correlation between even-movie IDMs from one range (rows) and odd-movie IDMs from another range (columns), normalized by the reliabilities of the corresponding ranges such that diagonal values equal 1. Diagonal values show same-range correlations that are consistently higher than off-diagonal values, demonstrating dimension-specific individual difference patterns. From left to right: early visual, ventral temporal, lateral visual, and posterior parietal-cingulate regions. (b) Same analysis after controlling for coarse-scale effects by partialing out voxelwise intersubject correlation (ISC) IDMs. The persistent diagonal structure indicates distinct, dimension-specific information in the residual, fine-grained patterns of individual differences. (c) Diagonal dominance (difference between on-diagonal correlation and mean off-diagonal correlation) remains significant across dimensional ranges before and after removing the coarse-grained effects. The diagonal dominance across ranks demonstrates that each dimensional range captures distinct aspects of individual neural processing during naturalistic viewing. Solid points show values from original matrices in a; lighter points show values after controlling for coarse-scale effects detected by conventional measures. All points shown are statistically significant (permutation tests, $p<0.05$).